# A meaningful prediction of functional decline in amyotrophic lateral sclerosis based on multi-event survival analysis

### v1

@@ Conclusion
This article proposes a novel and interpretable method for estimating the time to functional decline in ALS patients. We formulate this task as a multi-event survival problem and validate our method in the PRO-ACT dataset by training five covariate-based survival models to predict significant functional impairment with respect to five common functions. The proposed method has the following advantages over state-of-the-art approaches: (1) It explicitly predicts the time until functional decline is likely to occur for an individual patient. (2) It gives superior time-to-event estimates over commonly-used population estimators, such as the Kaplan-Meier. (3) It allows for individual counterfactual predictions, where certain covariates can be changed to see their effect on the predicted outcome. In conclusion, we recommend using the proposed method on patient records in clinical settings, subject to additional rigorous validation, to assess the risk of imminent functional decline and enable more personalized treatment planning.
@@ Acknowledgments
[…]
R.G was funded in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Alberta Machine Intelligence Institute (Amii), and the Canadian Institute for Advanced Research (CIFAR).
Authors Contributions
C.M.L conducted the data investigation and preparation, trained the models, performed the experiments, and wrote the manuscript. S.K and R.G conceived and initiated the research, provided supervision, and edited and prepared the manuscript for publication. The members of the Pooled Resource Open-Access ALS Clinical Trials Database (PRO-ACT) Consortium collected the PRO-ACT dataset.
Conflicts of Interest
All authors declare that they have no conflicts of interest.
Data Availability Statement
The PRO-ACT dataset can be downloaded from the Neurological Clinical Research Institute’s website (https://ncri1.partners.org/ProACT). Users must create an account to access the data, but account creation is free of charge.

### v2

@@ Conclusion
This article proposes a novel and interpretable method for estimating the time to functional decline in ALS patients. We formulate this task as a multi-event survival problem and validate our method in the PRO-ACT dataset by training five covariate-based survival models to predict significant functional impairment with respect to five common functions. The proposed method has the following advantages over state-of-the-art approaches: (1) It explicitly predicts the time until functional decline is likely to occur for an individual patient. (2) It gives superior time-to-event estimates over commonly used population estimators, such as the KM estimator. (3) It allows for individual counterfactual predictions, where certain covariates can be changed to see their effect on the predicted outcome. In conclusion, we recommend using the proposed method on patient records in clinical settings, subject to additional rigorous validation, to assess the risk of imminent functional decline and enable more personalized treatment planning.
@@ Acknowledgments
Data used in the preparation of this article were obtained from the Pooled Resource Open-Access ALS Clinical Trials (PRO-ACT) Database. The following organizations and individuals within the PRO-ACT Consortium contributed to the design and implementation of the PRO-ACT Database and/or provided data, but did not participate in the analysis of the data or the writing of this report: ALS Therapy Alliance, Cytokinetics, Inc., Amylyx Pharmaceuticals, Inc., Knopp Biosciences, Neuraltus Pharmaceuticals, Inc., Neurological Clinical Research Institute (MGH), Northeast ALS Consortium, Novartis, Orion Corporation, Prize4Life Israel, Regeneron Pharmaceuticals, Inc., Sanofi, Teva Pharmaceutical Industries, Ltd., and The ALS Association.